# Enhancing nanocrystal superlattice self-assembly near a metastable liquid binodal

### v1

@@ Abstract
Bottom-up assembly of nanocrystals (NCs) into ordered arrays, or superlattices (SLs), is a promising route to design materials with new functionalities, but the degree of control over assembly into functional structures remains challenging. Using electrostatics, rather than density, to tune the interactions between semiconductor NCs, we watch self-assembly proceeding through a metastable liquid phase. We systematically investigate the phase behavior as a function of quench conditions in situ and in real time using small angle X-ray scattering (SAXS). Through quantitative fitting to colloid, liquid, and SL models, we extract the time evolution of each phase and the system phase diagram, which we find to be consistent with short-range attractive interactions. Using the phase diagram’s predictive power, we establish control of the self-assembly rate over three orders of magnitude, and identify one- and two-step self-assembly regimes, with only the latter implicating the metastable liquid as an intermediate. Importantly, the presence of the metastable liquid increases SL formation rates relative to the equivalent one-step pathway, and SL order counterintuitively increases with the rate, revealing a highly desirable and generalizable kinetic strategy to promote and enhance ordered assembly.
@@ 0.0.1 Keywords:
The range of interparticle interactions relative to the size of the interacting particles determines the qualitative features of a system’s equilibrium phase diagram1, 2, 3. Atomic systems have a long range of the interactions relative to atomic scales, yielding temperature-density phase diagrams like in Figure 1a4, 5. As this range is decreased, liquid phase stability diminishes and the gas-liquid binodal or, for a colloidal system, the colloid-liquid binodal (blue) shifts to lower temperature. At a range of less than $\sim$20of the particle diameter2, the colloid-liquid binodal sits entirely below the colloid-solid binodal (Figure 1b). Consequently, as an alternative to typical one-step crystallization directly from the colloidal phase (Figure 1c top), systems like proteins with short range interactions may crystallize via a metastable liquid intermediate (Figure 1c bottom)1, 6, 3, 7, 8, 9, 10. For proteins and in simulations1, 6, 3, 11, 8, 12, 10, the two-step crystallization pathway has been shown to speed up crystallization over the equivalent one-step rate without increasing the crystallization driving force since the high density of particles in the liquid phase facilitates the order fluctuations necessary to form the crystal. Many questions, however, remain about the generality and underlying mechanisms of two-step crystallization. In particular, as seen in colloid-polymer mixtures and nanocrystal (NC) systems with similar interaction ranges to many protein systems, gelation often occurs before a metastable liquid can form. How subtle differences in the interparticle interactions alter crystallization outcomes is therefore essential to elucidate so that design principles can leverage the metastable liquid available via short-range interactions13, 14, 15, 10. Recently, colloid-metastable liquid phase separation as a precursor to crystallization has been seen in electrostatically stabilized PbS semiconductor NCs dispersed in electrolyte solutions 16. To fully uncover the crystallization mechanism and the role of underlying interparticle interactions requires systematic characterization of self-assembly trajectories in many different regimes.
@@
Measuring the phase evolution of NC systems can be challenging due to the fast time- and small length-scales intrinsic to NCs. Typically, the equilibrium phase behavior of colloids is measured using optical microscopy and scattering methods17, 18, 19, 20, yet optical methods insufficiently capture sub-diffraction nanoscale evolution. Recent small angle X-ray scattering (SAXS) has measured the evolving phase coexistence, in situ and in real-time, as colloidal NCs convert to superlattice (SL),21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32 an ordered solid of NCs33, 34, 35, 36, 37. Compared to visible light, X-rays provide more detailed, nanoscale information that is especially useful to distinguish different phases that do not separate on a macroscopic length scale on self-assembly time scales. Distinguishing coexisting colloidal, liquid, and solid phases remains challenging, however, due to their overlapping features in reciprocal space, limiting qualitative deductions from scattering patterns and common quantitative Bragg peak analyses38. Special care is therefore needed to accurately extract the phase behavior of NC systems as well as the kinetics of their phase transformations from SAXS data.
@@
Despite the fast time- and short length-scales of NC phase behavior, as well as the challenge of distinguishing phases in SAXS data, we systematically characterize the self-assembly of SLs from electrostatically stabilized PbS NCs using in situ SAXS, capturing multiple kinetic pathways, including the evolution of the liquid-state intermediate. To do so, we leverage the gas-tight reactor that we recently developed to study Au NC SL self-assembly21. We developed a model to fit SAXS patterns that quantitatively distinguishes between colloidal, liquid, and SL phases. From these fits we obtain the evolution of each phase as a function of time and quench depth ($k_{\mathrm{B}}T/u_{0}$, where $u_{0}$ is the depth of the interparticle potential) to deduce self-assembly mechanisms and kinetic rates that are not retrievable ex situ. By electrostatically tuning interparticle interactions, we reliably control self-assembly rates over three orders of magnitude as well as the mechanism: crystallization of SLs occurs either directly from the colloidal phase or in a two-step process with a metastable liquid intermediate. Our results suggest that the two-step pathway increases crystallization rates over the equivalent one-step pathway while simultaneously increasing SL order, a highly attractive kinetic strategy. Through comparison of the PbS NC system to its Au NC counterpart,21 we clarify how specific chemical and physical properties impact interparticle interactions and self-assembly, identifying the balance of interactions between the PbS NCs that avails the liquid phase and control of crystallization. In addition to establishing these design principles, our experimental and analysis methods should reveal the kinetic and thermodynamic complexity of a wide range of other nanosystems at the forefront of functional materials design33, 34, 35, 36, 37.
@@
Figure 1: Impact of interparticle potentials on thermodynamics and kinetics of self-assembly. a. Effective temperature ($k_{\mathrm{B}}T/u_{0}$) vs. volume fraction ($\phi$) phase diagram (top) and interaction potential vs. interparticle separation (bottom) for particles interacting via long-range interparticle potentials. b. Effective temperature vs. volume fraction phase diagram (top) and interaction potential vs. interparticle separation (bottom) for particles interacting via short-range interparticle potentials. In both a and b, $u_{0}$ is the magnitude of the minimum value of the interaction potential. c. Schematic of one-step (top) and two-step (bottom) SL formation pathways.
@@
Phases and phase evolution of self-assembling short-range-interacting nanocrystals. To determine the effect of a metastable liquid phase on the self-assembly of NC SLs, we study 5.8$\pm$0.3 nm diameter electrostatically stabilized PbS NCs (Figure S1a-c) dispersed in a polar solvent (N-methylformamide (NMF) mixed with N,N-dimethylformamide (DMF)), with multivalent, inorganic thiostannate ($\mathrm{Sn}_{2}\mathrm{S}_{6}^{4-}$) ligands on their surface16. We quench to a condensed phase (Figure S1d) by injecting a solution of additional multivalent salt ($\mathrm{K}_{3}\mathrm{AsS}_{4}$) into NC solutions of varying volume fraction. The salt screens the electrostatic repulsion to achieve attractive interactions between NCs. Over tens of single-shot experiments we systematically tune the solution ionic strength (see Methods) from 0.9 to 1.5 M to vary the quench depth, i.e., driving force for self-assembly, and we tune the volume fraction of NCs from $\phi_{0}\sim 3\times 10^{-5}$ to $8\times 10^{-3}$. We thus traverse a two-dimensional (2D) parameter space like those in the phase diagrams in Figures 1a,b.
@@
To follow the entire self-assembly process in situ, we employ a gas-tight reactor previously developed for SAXS21. A quartz cuvette with thin, X-ray-transparent windows initially contains NCs colloidally dispersed in an NMF/DMF mixture (Figure S2). The cuvette is connected via tubing to a syringe on a syringe pump that injects a controlled amount of $\mathrm{K}_{3}\mathrm{AsS}_{4}$ in NMF, inducing self-assembly. A magnetic stir bar in the cuvette ensures the X-rays sample the solution representatively. We collect 2D SAXS detector images before, during, and after salt injection with down-to-millisecond time resolution and for up to two hours post-injection. By azimuthally averaging detector images, we obtain one-dimensional SAXS patterns, $I(q)$, the scattered intensity as a function of momentum transfer, $q$. The time-evolution of $I(q)$ reveals the detailed evolution of the system’s coexisting phases, including the liquid phase density and SL crystallinity.
@@
Figure 2a shows an example of a time lapse with key time points over the course of an hour. Before salt injection the SAXS patterns show the decaying oscillatory scattering characteristic of 5.8 nm colloidal NCs (top curve). After the salt injection, the first five Bragg peaks of an fcc SL appear, e.g., (111) at $q\sim$0.11 Å1superscriptÅ$^{-1}$, (200) at 0.12 Å1superscriptÅ$^{-1}$, growing progressively until the system approaches equilibrium. Additional examples with distinct combinations of NC volume fractions and quench depths are found in Figure S3 and Figure S4. Globally, Bragg peaks emerge more quickly at deeper quenches and at higher NC volume fractions, as expected. The Bragg peaks are broader at deeper quenches, indicating that at deeper quenches the SL domains are smaller and more disordered.38. While the colloidal form factor and SL Bragg peaks are discernible by eye, noting the absence or presence of a liquid phase is not.
@@
To adequately describe the SAXS patterns over a range of NC concentrations and ionic strengths, we expanded a model that we previously developed21 to fit SAXS patterns of systems containing colloidal and SL phases in coexistence by including an additional term for the scattering from a NC liquid phase. We model the background-subtracted scattered intensity as $I(q)=I_{\mathrm{colloid}}(q)+I_{\mathrm{liquid}}(q)+I_{\mathrm{SL}}(q)+I_{\mathrm{LQ}}(q)$ (see Methods, Figures S5, S6a-c and associated text). Here, $I_{\mathrm{colloid}}(q)$ is the scattered intensity from dilute, polydisperse hard spheres, $I_{\mathrm{liquid}}(q)$ is the scattered intensity from a liquid of hard spheres at a volume fraction $\phi_{\mathrm{liquid}}$, $I_{\mathrm{SL}}(q)$ is the scattered intensity from finite-sized fcc SLs, and $I_{\mathrm{LQ}}(q)$ is the scattered intensity at low $q$ due to the finite size of high-density structures formed. The model assumes that the different components are statistically uncorrelated, which we validated on simulated systems (see SI, Figure S6d,e). We find that the model fits the experimental data well at all time points, NC volume fractions, and quench depths and also provides detailed, instantaneous information about each phase, such as the SL crystallinity and liquid phase density. Examples are shown near equilibrium in Figure 2b and Figure S7.
@@
Depending on the NC volume fraction and quench depth, we find cases in which the system contains strictly colloid and SL in coexistence (Figure S7a) or cases for which the colloid coexists with the SL and liquid phases (Figures 2b, S7b-d). To consolidate all of the long-time equilibrium behavior of the system as a function of the NC volume fraction and quench depth parameters explored, we indicate experimentally-determined points on the binodal curves that delineate the phase boundaries of this system’s phase diagram in Figure 2c (see Methods). These points specify the volume fraction, or density, of NCs in the colloidal or SL phases as a function of quench depth within the binodal curves. Even though the NC solvation has characteristics that go beyond a simple double layer,39, 40, 41 we use the Debye length, $\lambda$, of the solution as a figure of merit for the quench depth on the vertical axis since it combines information about the varying ionic strengths of the solution and the (fixed) dielectric constant of the solvent. Although this choice does not incorporate the impact of the steric and van der Waals contributions from the NCs into the quench depth parameter, these forces should not vary from quench to quench since the same NC stock solution was used for all measurements.
@@
To determine the points at each quench depth on the low-density side of the phase diagram in Figure 2c, i.e., the volume fraction of the colloidal phase, $\phi_{\mathrm{colloid}}$, we multiply the fraction of NCs remaining in the colloidal phase post-quench by the total volume fraction of NCs in the system. We calculate the high-density colloid-SL binodal points’ locations from the corresponding densities of the SL phase, $\phi_{\mathrm{SL}}$, at each quench depth based on the fcc (111) Bragg peak positions. Here, we include an additional colloid-metastable liquid binodal (grey data points and curve in Figure 2c) that bounds the colloid-metastable liquid coexistence. The high-density side of this binodal is defined by the liquid volume fraction, $\phi_{\mathrm{liquid}}$, extracted from the $I_{\mathrm{liquid}}(q)$ component of the fits. The low-density side is obtained from the density of the colloidal phase in coexistence with the metastable liquid, which is extracted from cases for which we observe a separation of time scales between liquid and SL formation (grey circle in Figure 2d). In order to resolve the binodals between $\phi$ $\sim$ 0.01 and 0.3, higher total NC volume fractions that are difficult to stabilize would be required. Nevertheless, experiments for which we injected insufficient salt to condense the colloid into either an ordered or disordered dense phase allow us to bound the colloid-SL binodal location at low volume fraction (open green squares in Figure 2c,d). Although the critical density is experimentally inaccessible, we constrain the location and height of the low-density colloid-metastable liquid binodal based on phase diagram locations that resulted in one-step SL formation pathways (open purple circles in Figure 2c,d). Another phase diagram obtained in the same way for a different stock solution of PbS NCs is shown in Figure S8.
@@
Figure 2: Quantitative analysis of quench- and volume fraction-dependent SAXS patterns. a. One-dimensional SAXS patterns as a function of time after injection of excess salt for an example experiment using NC volume fraction $\phi_{0}=2.8\times 10^{-3}$ and a solution ionic strength (quench depth) of 1.05 M. Additional examples at other volume fractions and quench depths can be found in Figure S3. Quantitative fit (red) of SAXS pattern obtained using colloidal NC (green), liquid (blue), and SL (yellow) phases at the same NC volume fraction and quench depth as in a after equilibrating $>$ 1 hour post quench for an experiment at $\phi_{0}=8.5\times 10^{-4}$ and a solution ionic strength of 1.2 M. Additional examples of fits corresponding to the four cases in Figure S3 can be found in Figure S7. c. Quantitative phase diagram for electrostatically stabilized PbS NCs obtained from experimental observations. Black data points fall on the colloid-SL binodal and grey data points fall on the colloid-liquid binodal. Vertical error bars indicate the standard deviations of the Debye lengths, $\lambda$, of the solutions based on the uncertainty of the volume and concentration of the injected salt solution. Horizontal error bars indicate the standard deviations of the colloidal, liquid, and SL volume fractions due to the same volume uncertainty and uncertainty from SAXS fitting. Black (colloid-SL) and grey (colloid-liquid) phase boundary curves are sketched as a visual guide. d. Close-up of low-density region of the phase diagram with additional points included. The open green squares indicate ($\phi,\lambda$) phase diagram locations in which the system is found only in the colloidal state. The open purple circles indicate ($\phi,\lambda$) locations in which no liquid component was required in the model to fit the SAXS patterns. The relative fractions of the different phases observed are represented for four distinct regimes whose fits are shown in Figure S7 as pie charts color coded to match the fits in a (colloidal NC: green, liquid: blue, SL: yellow).
@@
Figure 3: Kinetic analysis of quench- and volume fraction-dependent SAXS. a-d. Fraction of NCs in the colloidal (green), liquid (blue), and SL (yellow) phases as a function of time for NC volume fractions and quench depth cases covered in Figure S3a-d and Figure S7a-d. e. Map of experimentally explored parameter space, indicating approximate progression of regions exhibiting the four kinetic behaviors shown in a-d. Black and grey curves sketch the approximate colloid-SL and colloid-metastable liquid binodals, respectively, as in Figure 2c. f. Distinct crystallization pathways illustrated on cartoon free energy surfaces as a function of NC order and density. Arrow widths indicate relative rates within each pathway. Blue shades indicate negative relative free energies; red shades indicate positive relative free energies.
@@
Deducing features of electrostatically tunable nanoscale interparticle potentials. Having described our SAXS data and analysis, we now discuss the implications of our findings in terms of inferred interaction potentials, advantages of SL formation pathways, and emerging design principles. Previously, the phase behavior of these NCs was determined qualitatively via observing macroscopic phase separation in cuvettes at long times post-quench16. Here, we go beyond this initial work by directly and quantitatively measuring the phase diagram of these NCs over a large range of volume fractions and quench depths (Figure 2c,d). Understanding the different contributions to the associated interparticle potentials and their dependence on external control parameters enables the fine tuning of interactions necessary to optimize ordered SL self-assembly for short-range interacting nanoscale particles. Both the equilibrium phase coexistence and kinetic mechanisms of SL self-assembly from electrostatically stabilized NCs are ultimately determined by the underlying depth and range of interactions between NCs. The range of interactions can be inferred from the extracted phase diagram, which is consistent with the phase diagrams of other particles interacting with short-range interparticle potentials. Based on the Noro-Frenkel law of corresponding states2 and the effective size of the NCs ($\sim$6.4-7.8 nm, see Methods), we infer the surface-to-surface range of the interparticle potential to be 1.9-3.6 nm in order for the liquid phase to be metastable. Simulation studies bound interaction well depths of short-range interacting systems to $2k_{\mathrm{B}}T<u_{0}<6k_{\mathrm{B}}T$ range3.

### v2

@@ Abstract
Achieving predictive control over crystallization using non-classical nucleation while avoiding kinetic traps is a roadblock toward designing materials with new functionalities. We address these challenges by inducing bottom-up assembly of nanocrystals (NCs) into ordered arrays, or superlattices (SLs). Using electrostatics, rather than density, to tune the interactions between particles, we watch self-assembly proceeding through a metastable liquid phase. We systematically investigate the phase behavior as a function of quench conditions in situ and in real time using small angle X-ray scattering (SAXS). Through quantitative fitting to colloid, liquid, and SL models, we extract the time evolution of each phase and the system phase diagram, which we find to be consistent with short-range attractive interactions. Using the phase diagram’s predictive power, we establish control of the self-assembly rate over three orders of magnitude, and identify one- and two-step self-assembly regimes, with only the latter implicating the metastable liquid as an intermediate. Importantly, the presence of the metastable liquid increases SL formation rates relative to the equivalent one-step pathway, and SL order counterintuitively increases with the rate, revealing a highly desirable and generalizable kinetic strategy to promote and enhance ordered assembly.
@@ 0.0.1 Keywords:
The range of interparticle interactions relative to the size of the interacting particles determines the qualitative features of a system’s equilibrium phase diagram1, 2, 3. Atomic systems have a long range of the interactions relative to atomic scales, yielding temperature-density phase diagrams like in Figure 1a4, 5. As this range is decreased, liquid phase stability diminishes and the gas-liquid binodal or, for a colloidal system, the colloid-liquid binodal (grey) shifts to lower temperature. The colloid-liquid binodal is a fluid-fluid binodal in which the colloidal phase is the low-density fluid and the liquid phase is the high-density fluid composed of densely packed colloidal particles. At a range of less than $\sim$20of the particle diameter2, the colloid-liquid binodal (Figure 1b, grey curve) sits entirely below the colloid-solid binodal (Figure 1b, black curve), rendering the liquid phase metastable 1, 2, 6. Consequently, as an alternative to typical one-step crystallization directly from the colloidal phase (Figure 1c top), systems like proteins with short range attractive interactions may crystallize via a metastable liquid intermediate (Figure 1c bottom)1, 7, 3, 8, 9, 10, 11. For proteins and in simulations1, 7, 3, 12, 9, 13, 11, the two-step crystallization pathway has been shown to speed up crystallization over the equivalent one-step rate without increasing the crystallization driving force since the high density of particles in the liquid phase facilitates the order fluctuations necessary to form the crystal. Many questions, however, remain about the generality and underlying mechanisms of two-step crystallization. In particular, as seen in colloid-polymer mixtures and nanocrystal (NC) systems with similar interaction ranges to many protein systems, gelation often occurs before a metastable liquid can form. How subtle differences in the interparticle interactions alter crystallization outcomes is therefore essential to elucidate so that design principles can leverage the metastable liquid available via short-range interactions14, 15, 16, 11. Recently, colloid-metastable liquid phase separation as a precursor to crystallization has been seen in electrostatically stabilized PbS semiconductor NCs dispersed in electrolyte solutions 17. To fully uncover the crystallization mechanism and the role of underlying interparticle interactions requires systematic characterization of self-assembly trajectories in many different regimes.
@@
Measuring the phase evolution of nanoscale systems can be challenging due to the fast time- and small length-scales intrinsic to them. Typically, the equilibrium phase behavior of colloids is measured using optical microscopy and scattering methods18, 19, 20, 21, yet optical methods insufficiently capture sub-diffraction nanoscale evolution. Recent small angle X-ray scattering (SAXS) has measured the evolving phase coexistence, in situ and in real-time, as colloidal NCs convert to superlattice (SL),22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33 an ordered solid of NCs34, 35, 36, 37, 38. Compared to visible light, X-rays provide more detailed, nanoscale information that is especially useful to distinguish different phases that do not separate on a macroscopic length scale on self-assembly time scales. Distinguishing coexisting colloidal, liquid, and solid phases remains challenging, however, due to their overlapping features in reciprocal space, limiting qualitative deductions from scattering patterns and common quantitative Bragg peak analyses39. Special care is therefore needed to accurately extract the phase behavior of NC systems as well as the kinetics of their phase transformations from SAXS data.
@@
Despite the fast time- and short length-scales of NC phase behavior, as well as the challenge of distinguishing phases in SAXS data, we systematically characterize the self-assembly of SLs from electrostatically stabilized PbS NCs using in situ SAXS, capturing multiple kinetic pathways, including the evolution of the liquid-state intermediate. To do so, we leverage the gas-tight reactor that we recently developed to study Au NC SL self-assembly22. We developed a model to fit SAXS patterns that quantitatively distinguishes between colloidal, liquid, and SL phases. From these fits we obtain the evolution of each phase as a function of time and quench depth ($k_{\mathrm{B}}T/u_{0}$, where $u_{0}$ is the depth of the interparticle potential) to deduce self-assembly mechanisms and kinetic rates that are not retrievable ex situ. By electrostatically tuning interparticle interactions, we reliably control self-assembly rates over three orders of magnitude as well as the mechanism: crystallization of SLs occurs either directly from the colloidal phase or in a two-step process with a metastable liquid intermediate. Our results suggest that the two-step pathway increases crystallization rates over the equivalent one-step pathway while simultaneously increasing SL order, a highly attractive kinetic strategy. Through comparison of the PbS NC system to its Au NC counterpart,22 we clarify how specific chemical and physical properties impact interparticle interactions and self-assembly, identifying the balance of interactions between the PbS NCs that avails the liquid phase and control of crystallization. In addition to establishing these design principles, our experimental and analysis methods should reveal the kinetic and thermodynamic complexity of a wide range of other nanosystems at the forefront of functional materials design and provide a strategy to achieve similar order in place of gelation with larger scale colloids.34, 35, 36, 37, 38, 3, 40, 41.
@@
Figure 1: Impact of interparticle potentials on thermodynamics and kinetics of self-assembly. a. Effective temperature ($k_{\mathrm{B}}T/u_{0}$) vs. volume fraction ($\phi$) phase diagram (top) and sketch of interaction potential vs. interparticle separation (bottom) for particles interacting via long-range interparticle potentials. b. Sketch of effective temperature vs. volume fraction phase diagram (top) and sketch of interaction potential vs. interparticle separation (bottom) for particles interacting via short-range interparticle potentials. In both a and b, $u_{0}$ is the magnitude of the minimum value of the interaction potential. c. Schematic of one-step (top) and two-step (bottom) SL formation pathways.
@@
Phases and phase evolution of self-assembling short-range-interacting nanocrystals. To determine the effect of a metastable liquid phase on the self-assembly of NC SLs, we study 5.8$\pm$0.3 nm diameter electrostatically stabilized PbS NCs (Figure S1a-c) dispersed in a polar solvent (N-methylformamide (NMF) mixed with N,N-dimethylformamide (DMF)), with multivalent, inorganic thiostannate ($\mathrm{Sn}_{2}\mathrm{S}_{6}^{4-}$) ligands on their surface17. We quench to a condensed phase (Figure S1d) by injecting a solution of additional multivalent salt ($\mathrm{K}_{3}\mathrm{AsS}_{4}$) into NC solutions of varying volume fraction. The salt screens the initial electrostatic repulsion generated by the charged thiostannate ligands to achieve attractive interactions between NCs. Over tens of single-shot experiments we systematically tune the solution ionic strength (see Methods) from 0.9 to 1.5 M to vary the quench depth, i.e., driving force for self-assembly, and we tune the volume fraction of NCs from $\phi_{0}\sim 3\times 10^{-5}$ to $8\times 10^{-3}$. We thus traverse a two-dimensional (2D) parameter space like those in the phase diagrams in Figures 1a,b.
@@
To follow the entire self-assembly process in situ, we employ a gas-tight reactor previously developed for SAXS22. A quartz cuvette with thin, X-ray-transparent windows initially contains NCs colloidally dispersed in an NMF/DMF mixture (Figure S2). The cuvette is connected via tubing to a syringe on a syringe pump that injects a controlled amount of $\mathrm{K}_{3}\mathrm{AsS}_{4}$ in NMF, inducing self-assembly. A magnetic stir bar in the cuvette ensures the X-rays sample the solution representatively. We collect 2D SAXS detector images before, during, and after salt injection with down-to-millisecond time resolution and for up to two hours post-injection. By azimuthally averaging detector images, we obtain one-dimensional SAXS patterns, $I(q)$, the scattered intensity as a function of momentum transfer, $q$. The time-evolution of $I(q)$ reveals the detailed evolution of the system’s coexisting phases, including the liquid phase density and SL crystallinity.
@@
Figure 2a shows an example of a time lapse with key time points over the course of an hour. Before salt injection the SAXS patterns show the decaying oscillatory scattering characteristic of 5.8 nm colloidal NCs (top curve). After the salt injection, the first five Bragg peaks of an fcc SL appear, e.g., (111) at $q\sim$0.11 Å1superscriptÅ$^{-1}$, (200) at 0.12 Å1superscriptÅ$^{-1}$, growing progressively until the system approaches equilibrium. Additional examples with distinct combinations of NC volume fractions and quench depths are found in Figure S3 and Figure S4. Globally, Bragg peaks emerge more quickly at deeper quenches and at higher NC volume fractions, as expected. The Bragg peaks are broader at deeper quenches, indicating that at deeper quenches the SL domains are smaller and more disordered.39. While the colloidal form factor and SL Bragg peaks are discernible by eye, noting the absence or presence of a liquid phase is not.
@@
To adequately describe the SAXS patterns over a range of NC concentrations and ionic strengths, we expanded a model that we previously developed22 to fit SAXS patterns of systems containing colloidal and SL phases in coexistence by including an additional term for the scattering from a NC liquid phase. We model the background-subtracted scattered intensity as $I(q)=I_{\mathrm{colloid}}(q)+I_{\mathrm{liquid}}(q)+I_{\mathrm{SL}}(q)+I_{\mathrm{LQ}}(q)$ (see Methods, Figures S5, S6a-c and associated text). Here, $I_{\mathrm{colloid}}(q)$ is the scattered intensity from dilute, polydisperse hard spheres, $I_{\mathrm{liquid}}(q)$ is the scattered intensity from a liquid of hard spheres at a volume fraction $\phi_{\mathrm{liquid}}$, $I_{\mathrm{SL}}(q)$ is the scattered intensity from finite-sized fcc SLs, and $I_{\mathrm{LQ}}(q)$ is the scattered intensity at low $q$ due to the finite size of high-density structures formed. The model assumes that the different components are statistically uncorrelated, which we validated on simulated systems (see SI, Figure S6). We find that the model fits the experimental data well (see Figures S7, S8 ) at all time points, NC volume fractions, and quench depths and also provides detailed, instantaneous information about each phase, such as the SL crystallinity and liquid phase density. Examples are shown near equilibrium in Figure 2b and Figure S9.
@@
Depending on the NC volume fraction and quench depth, we find cases in which the system contains strictly colloid and SL in coexistence (Figure S9a) or cases for which the colloid coexists with the SL and liquid phases (Figures 2b, S9b-d). To consolidate all of the long-time equilibrium behavior of the system as a function of the NC volume fraction and quench depth parameters explored, we indicate experimentally-determined points on the binodal curves that delineate the phase boundaries of this system’s phase diagram in Figure 2c (see Methods). These points specify the volume fraction, or density, of NCs in the colloidal or SL phases as a function of quench depth within the binodal curves. We cannot explicitly know a potential well depth for the yet-unspecified short-range potential of the quenched NCs, so even though the NC solvation has characteristics that go beyond a simple double layer,42, 43, 44 we use the Debye length, $\lambda$, of the solution as a figure of merit for the quench depth on the vertical axis. The Debye length combines information about the varying ionic strengths of the solution and the (fixed) dielectric constant of the solvent. Although this choice does not incorporate the impact of the steric and van der Waals contributions from the NCs into the quench depth parameter, these forces should not vary from quench to quench since the same NC stock solution was used for all measurements. For clarity, the ionic strength is also indicated on the right-hand side of the plot.
@@
To determine the points at each quench depth on the low-density side of the phase diagram in Figure 2c, i.e., the volume fraction of the colloidal phase, $\phi_{\mathrm{colloid}}$, we multiply the fraction of NCs remaining in the colloidal phase post-quench by the total volume fraction of NCs in the system. We calculate the high-density colloid-SL binodal points’ locations from the corresponding densities of the SL phase, $\phi_{\mathrm{SL}}$, at each quench depth based on the fcc (111) Bragg peak positions. The volume fraction of the SL phase is lower than expected for typical fcc crystals due to the presence of ligand and electrolyte molecules in the interstices of the lattice, neither of which is included in the computed NC volume. Here, we include an additional colloid-metastable liquid binodal (grey data points and curve in Figure 2c) that bounds the colloid-metastable liquid coexistence. The high-density side of this binodal is defined by the liquid volume fraction, $\phi_{\mathrm{liquid}}$, extracted from the $I_{\mathrm{liquid}}(q)$ component of the fits. The low-density side is obtained from the density of the colloidal phase in coexistence with the metastable liquid, which is extracted from cases for which we observe a separation of time scales between liquid and SL formation (grey circle in Figure 2d). Limitations on our estimates of these binodals are discussion in the SI. In order to resolve the binodals between $\phi$ $\sim$ 0.01 and 0.3, higher total NC volume fractions that are difficult to stabilize would be required. We did not do so in this work due to the greater difficulty stabilizing the colloidal phase at high NC volume fractions. Nevertheless, experiments for which we injected insufficient salt to condense the colloid into either an ordered or disordered dense phase allow us to bound the colloid-SL binodal location at low volume fraction (open green squares in Figure 2d). Although the critical density is experimentally inaccessible, we constrain the location and height of the low-density colloid-metastable liquid binodal based on phase diagram locations that resulted in one-step SL formation pathways (open purple circles in Figure 2d). Another phase diagram obtained in the same way for a different stock solution of PbS NCs is shown in Figure S10.
@@
Figure 2: Quantitative analysis of quench- and volume fraction-dependent SAXS patterns. a. One-dimensional SAXS patterns as a function of time after injection of excess salt for an example experiment using NC volume fraction $\phi_{0}=2.8\times 10^{-3}$ and a solution ionic strength (quench depth) of 1.05 M. Additional examples at other volume fractions and quench depths can be found in Figure S3. b. Quantitative fit (red) of SAXS pattern obtained using colloidal NC (green), liquid (blue), and SL (yellow) phases at the same NC volume fraction and quench depth as in a after equilibrating $>$ 1 hour post quench for an experiment at $\phi_{0}=8.5\times 10^{-4}$ and a solution ionic strength of 1.2 M. Additional examples of fits corresponding to the four cases in Figure S3 can be found in Figure S7. c. Quantitative phase diagram for electrostatically stabilized PbS NCs obtained from experimental observations. Black points fall on the colloid-SL binodal and grey points fall on the colloid-liquid binodal. We sketched the curves between ϕsimilar-toitalic-$\phi\sim$ 0.01 and ϕsimilar-toitalic-$\phi\sim$ 0.3 with dashing as an interpolation between the solid curves dictated by experimental data; the dashed curves are not meant to be quantitative. Vertical error bars indicate the standard deviations of the Debye lengths, $\lambda$, of the solutions based on the uncertainty of the volume and concentration of the injected salt solution. Horizontal error bars indicate the standard deviations of the colloidal, liquid, and SL volume fractions due to the same volume uncertainty and uncertainty from SAXS fitting. Black (colloid-SL) and grey (colloid-liquid) phase boundary curves are sketched as a visual guide. d. Close-up of low-density region of the phase diagram with additional points included. The open green squares indicate ($\phi,\lambda$) phase diagram locations in which the system is found only in the colloidal state. The open purple circles indicate ($\phi,\lambda$) locations in which no liquid component was required in the model to fit the SAXS patterns. The relative fractions of the different phases observed are represented for four distinct regimes whose fits are shown in Figure S7 as pie charts color coded to match the fits in a (colloidal NC: green, liquid: blue, SL:yellow).
@@
Figure 3: Kinetic analysis of quench- and volume fraction-dependent SAXS. a-d. Fraction of NCs in the colloidal (green), liquid (blue), and SL (yellow) phases as a function of time for NC volume fractions and quench depth cases covered in Figure S3a-d and Figure S9a-d (see SI for further details). (The absence of data between $\sim$10 and $\sim$19 min in panel d is due to a pause incurred when changing detector sampling frequencies but does not compromise the interpretation of the kinetics for the quench conditions probed.) e. Map of experimentally explored parameter space, indicating approximate progression of regions exhibiting the four kinetic behaviors shown in a-d. Black and grey curves sketch the approximate colloid-SL and colloid-metastable liquid binodals, respectively, as in Figure 2c. f. Distinct crystallization pathways illustrated on cartoon free energy surfaces as a function of NC order and density. Arrow widths indicate relative rates within each pathway. Blue shades indicate negative relative free energies; red shades indicate positive relative free energies.
@@
Deducing features of electrostatically tunable nanoscale interparticle potentials. Having described our SAXS data and analysis, we now discuss the implications of our findings in terms of inferred interaction potentials, advantages of SL formation pathways, and emerging design principles. Previously, the phase behavior of these NCs was determined qualitatively via observing macroscopic phase separation in cuvettes at long times post-quench17. Here, we go beyond this initial work by directly and quantitatively measuring the phase diagram of these NCs over a large range of volume fractions and quench depths (Figure 2c,d). Understanding the different contributions to the associated interparticle potentials and their dependence on external control parameters enables the fine tuning of interactions necessary to optimize ordered SL self-assembly for short-range interacting nanoscale particles. Both the equilibrium phase coexistence and kinetic mechanisms of SL self-assembly from electrostatically stabilized NCs are ultimately determined by the underlying depth and range of interactions between NCs. The range of interactions can be inferred from the extracted phase diagram, which is consistent with the phase diagrams of other particles interacting with short-range interparticle potentials. Based on the Noro-Frenkel law of corresponding states2 and the effective size of the NCs ($\sim$6.4-7.8 nm, see Methods), we infer the surface-to-surface range of the interparticle potential to be 1.3-1.6 nm in order for the liquid phase to be metastable. Simulation studies bound interaction well depths of short-range interacting systems to $2k_{\mathrm{B}}T<u_{0}<6k_{\mathrm{B}}T$ range3.